OpenStudy (anonymous):

I need help.....if x,y>0, log y(x)+(x) log x(y)=10/3 and xy=144 then what is the value of x+y/2? homework problem please...

OpenStudy (anonymous):

if x,y>0, log y(x)+(x) log x(y)=10/3 and xy=144 then what is the value of x+y/2?

OpenStudy (anonymous):

Please help me. I dont even know where to start......

OpenStudy (anonymous):

Log ofy(x) +x = what?

OpenStudy (anonymous):

use x=144/y or y=144/x to get the equation into one variable

OpenStudy (anonymous):

if x,y >0, log of y(x) +(x)=10/3 log of x(y)=10/3 and xy=144 then what is the value of x+y/2? Does that answer your question?

OpenStudy (anonymous):

and remember the rules for rewriting logs

OpenStudy (anonymous):

am i supposed to substitute for all the variables?

OpenStudy (anonymous):

and should I substitute for the bases as well?

OpenStudy (anonymous):

write it out more clearly, one equation per line using parenthesis to show what is in the log

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x,y>0 log y (x)+log x(y)=10/3

OpenStudy (anonymous):

and xy=144 what is the value of (x+y)/2?

OpenStudy (anonymous):

when you write out the log you need to show what is part of the log: log(x)y is different than log(xy) and so on

OpenStudy (anonymous):

put the parenthesis first after to the log indicating that what is in the parenthesis is part of the log, that should be how you write it on your homework so it's clear

OpenStudy (anonymous):

log base y (x)+ log base x (y)

OpenStudy (anonymous):

Do you mean like what he just wrote?^

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Not you Jonathan. I'm talking to the people tryin to help.....

OpenStudy (anonymous):

Try reposting the problem properly because I dont think I am posting it right.